OpenStudy (kendricklamar2014):

A box contains 95 pink rubber bands and 90 brown rubber bands. You select a rubber bands at random from the box. Find each probability. Write the probability as a fraction in simplest form. A. Find the theoretical probability of selecting a Pink rubber band. B. Find the theoretical probability of selecting a Brown rubber band. C. You repeatedly choose a rubber band from the box, record the color, and put the rubber band back in the box. The results are shown in the table below. Find the experimental probability of each color based on the table.

OpenStudy (kendricklamar2014):

Outcome Occurrence Pink 36 Brown 33

OpenStudy (texaschic101):

A. 95/185 reduces to 19/37 (probability for pink) B. 90/185 reduces to 18/37 (probability for brown) I do not know the other one. I am not the best at probabilities, so you might better get somebody to check and see if I did this correctly

OpenStudy (isaiah.feynman):

@texaschic101 you are correct.

OpenStudy (texaschic101):

thanks for the confirmation...I am not good at these, so thanks for checking :)

OpenStudy (kendricklamar2014):

Can u help me with C

ganeshie8 (ganeshie8):

``` Outcome Occurrence Pink 36 Brown 33 ```

ganeshie8 (ganeshie8):

total trials = 36 + 33 = 99

ganeshie8 (ganeshie8):

P(Pink) = 36/99 = ? P(Brown) = 33/99 = ?

ganeshie8 (ganeshie8):

^^ thats for part C

OpenStudy (kendricklamar2014):

ok

OpenStudy (kendricklamar2014):

36/99=0.36 33/99=0.33

OpenStudy (kendricklamar2014):

@ganeshie8

ganeshie8 (ganeshie8):

obviously they're wrong, the reason is : they dont add up to 1

ganeshie8 (ganeshie8):

I made a mistake earlier while adding : 36 + 33 = 69 (NOT 99)

ganeshie8 (ganeshie8):

work it again

ganeshie8 (ganeshie8):

P(Pink) = 36/69= ? P(Brown) = 33/69 = ?

OpenStudy (kendricklamar2014):

36/69=0.52 33/69= 0.47

OpenStudy (kendricklamar2014):

@ganeshie8

ganeshie8 (ganeshie8):

36/69=0.52 33/69= 0.48 look good

OpenStudy (kendricklamar2014):

is that it?

OpenStudy (kendricklamar2014):

@ganeshie8

ganeshie8 (ganeshie8):

thats it !

OpenStudy (kendricklamar2014):

thanks, can u help me with one more please

OpenStudy (kendricklamar2014):

@ganeshie8